# PoliPrompt: A High-Performance Cost-Effective LLM-Based Text Classification Framework for Political Science

### v1

@@ 4.1 Classifying BBC News Reports Topics
Figure 3: Analyzing BBC Topics: Comparison of F1 Scores Across Different Methods
4.2 Measuring Public Opinion Toward Brett Kavanaugh’s SCOTUS Nomination
In the second experiment, we focus on demonstrating the effectiveness of our second stage—the prompt generator—in accurately labeling nuanced political science concepts.
In their 2023 paper, Bestvater and Monroe argue that sentiment and stance are fundamentally distinct concepts, a difference that traditional sentiment analysis tools like the VADER dictionary and the Lexicoder Sentiment Dictionary (LSD) often fail to capture. While sentiment analysis gauges the emotional tone of a document—whether positive, negative, or neutral—stance identification determines the author’s position on a specific issue, which may not align directly with the sentiment.
Bestvater and Monroe explore this distinction by analyzing a corpus of tweets about Kavanaugh’s nomination(2023). Their study compares the effectiveness of various text classifiers in identifying stance and concludes that, for many stance detection tasks, training a new supervised classifier on a hand-labeled dataset yields more accurate results than relying on existing models or dictionaries designed for sentiment analysis. The key findings from their study are summarized in the Appendix Table LABEL:tab:metrics_(Bestvater_and_Monroe).
To assess the ability of our method to distinguish nuanced political science concepts and generate accurate predictions, we applied it to the human-labeled dataset from Bestvater and Monroe (2023), which serves as the ground truth benchmark. We adopted the same instructions given to the human coders in their study and used them as the foundational heuristic prompt for two weaker language models (Mistral-medium and GPT-3.5). The simple heuristic prompt is detailed in Figure 4:
@@ 4.2 Measuring Public Opinion Toward Brett Kavanaugh’s SCOTUS Nomination
Using this 0-shot prompt, both Mistral-medium and GPT-3.5-turbo achieved F1 scores below 60%, as the weak LLMs struggled to differentiate between sentiment and stance. For instance, these LLMs mistakenly labeled the following Tweet as ”oppose”:
@@
RT @atensnut Democrats can’t just “move on” and jump on the bandwagon of Sketchy allegations against Kavanaugh, without accepting the egregiousness of turning their backs on the victims of Bill Clinton.
@@
The LLMs categorized this as ”oppose” because the tweet conveys a tone of negativity and distrust, implicitly criticizing the Democratic Party for supporting allegations against Kavanaugh while highlighting their perceived hypocrisy. This example illustrates how weak LLMs, when using a simple heuristic prompt, can easily confuse sentiment with stance.
@@
We then provided weak models with enhanced prompts produced by LLM in the second stage of our method, shown in Figure 4, offering a clear and concise guideline that significantly refined the LLMs’ judgment. Notably, the enhanced prompt goes beyond mere factual classification; for instance, it correctly identifies tweets as ”approve” when they present evidence that undermines Kavanaugh’s accusers. Additionally, the enhanced prompt effectively captures and summarizes emotional tones that could be interpreted as supportive of Kavanaugh, such as mocking his opponents and expressing frustration toward actions obstructing his confirmation. These nuances are easily confused with opposition when using a simple heuristic prompt, as the negative tone and sentiment of such tweets can misleadingly suggest an opposing stance. This simple guideline ensures that both factual content and emotional cues are considered, leading to more accurate and nuanced classifications. The enhanced prompt resulted in substantial improvements, with Mistral-medium and GPT-3.5 achieving impressive F1 scores of 91.69% and 92.59%, respectively. These results mark a significant improvement over the initial performance achieved with the simple heuristic prompt, showing approximately a 36% increase in F1 score for zero-shot prompting, as illustrated in Figure 5. Similarly, in the four-shot [^7] Detailed few-shot prompts are provided in Appendix LABEL:fig:Kavanaugh_few-shot_prompt. prompting scenario, our enhanced prompt enabled the weak LLMs to achieve over a 20% improvement, demonstrating the importance of having an enhanced prompt in generating the right classifications and label [^8] The detailed resulting performance metrics for 4-shot promptings are presented in Appendix LABEL:tab:Kavanaugh-4-shot-metrics. Essentially, by incorporating the final stage of chain-of-thought reasoning and the “Mistral-large-latest” judge model, our approach further enhanced the F1 scores, surpassing traditional supervised classifiers and reaching impressive levels above 95%..
Our method demonstrates superior performance compared to raw LLM predictions, largely due to the enhanced prompt generated in the second stage of our approach. This underscores an important consideration for political scientists: when using LLMs for text classification, relying on raw labeling alone can lead to significant misclassifications, as LLMs may struggle with nuanced concepts like sentiment and stance. However, by providing a few exemplars and instructing the LLM to generate a more refined prompt with simple, targeted guidelines, we achieve outstanding results. Furthermore, our method outperforms both dictionary-based and traditional supervised learning approaches, all without the need to train a new model—contrary to the recommendation by Bestvater and Monroe (2023) for improved classification. We successfully minimized the cost associated with model training while maximizing classification accuracy.
@@
Figure 4: 0-Shot Prompts for Analyzing Tweets about Brett Kavanaugh
@@
Figure 5: Measuring Opinion toward Kavanaugh: Comparison of F1 Scores Across Different Methods
@@ 4.3 Classifying Campaign Ads Tones in the 2018 Election
The medium through which political communication is delivered plays a critical role in shaping the message’s tone and its audience reach. In a recent study, Fowler and her colleagues examined the impact of Facebook as a medium on the tone of political advertisements (2021). They proposed that ads on Facebook are more likely to adopt a negative tone compared to other platforms.
@@
We applied the same three-stage approach as in our previous two experiments, this time focusing solely on reporting the final post-process results of the Chain-of-Thought (CoT) and Judge prompting methods. We evaluated their performance under both zero-shot and few-shot prompting conditions [^10] Detailed raw and in-process results are available in the Appendix.. The final F1 scores, as shown in Table 2, indicate notably low performance in the ”attack” and ”contrast” categories. Specifically, the average F1 score for ”attack” is approximately 0.55, while for ”contrast,” it is even lower, averaging around 0.5. It is important to note that these F1 scores were calculated against gold-standard human labels, which are particularly noisy in this dataset, potentially contributing to the reduced accuracy.
@@
Class CoT (k=0) Judge (k=0) CoT (k=6) Judge (k=6)
promote 0.9285 0.9323 0.9260 0.9314
contrast 0.5385 0.4567 0.5180 0.4978
attack 0.5719 0.5368 0.5648 0.5856
@@
These findings facilitated a re-examination of the data, particularly in cases where LLM-generated labels differ, as these discrepancies may indicate conceptual confusion and could lead to controversial or inconsistent decisions. Examples of such label inconsistencies, as shown in Table 3, underscore the need to question the consistency and reliability of the so-called gold standard human labels. In the first text, which was labeled as ”contrast” by human coders, the primary focus is on criticizing career politicians for failing to prevent the destruction of industries, without directly contrasting specific policies with those proposed by others. On the contrary, the second piece, labeled as ”promote” by human coders, explicitly criticizes the Republican Party in the state legislature while calling for support for other candidates. This message could reasonably be interpreted as both a promotion of alternatives and a contrast with the criticized party. Despite the similarities between these two ads—both of which criticize opponents and call for the support of others—the human labels differed. On the other hand, GPT and the final chain-of-thought method remained consistent in their classifications, demonstrating a more stable and coherent approach to labeling.
@@
Notes: The left panel of this figure shows the effect of medium on campaign ad tones across different labeling approaches, including initial human-labeled data before removing controversial labels. The right panel presents regression estimates after excluding the controversial observations.
Given the noise within the human-labeled data, we proceeded to remove all controversial observations – the observations are labeled differently from two weak LLMs, and Table 4 highlights the substantial improvements in F1 scores across all categories following this adjustment. Notably, the ”contrast” and ”attack” categories saw over a 30% increase in F1 scores. After this refinement, we reran the downstream fixed effects model. The right panel of Figure 6 shows that the estimates obtained through automated classification now closely align with those derived from human-labeled data [^12] The right panel displays identical estimates for both the chain-of-thought and judge models because mismatched observations were removed. As a result, the remaining labels for both models are consistent, leading to the same regression estimates..
@@
Additionally, Table 2 revealed that zero-shot metrics surpassed few-shot performance. This anomaly can be attributed to the sampling of 80 examples for few-shot prompting during the preprocessing stage. Given the noise within the human labels, these samples might have included incorrectly labeled examples, impacting the accuracy of few-shot promptings. This reliance on potentially flawed data could degrade the performance of few-shot prompts to levels below that of zero-shot prompts, which remain uninfluenced by human input.
This experiment underscores an additional advantage of our method: its ability to identify and flag problematic human labels. While human coding has traditionally been regarded as the gold standard, this assumption falters when dealing with controversial classifications [^13] The inconsistency among human coders is well-documented in the literature.. Our findings urge caution against placing absolute trust in human coding, particularly when relying on a single coder, as disagreements are inevitable.
@@
Table 4: Analyzing Ads Tones: F1 Scores after Addressing Discrepancies: CoT Prompting vs. Judge Prompting (0-Shot)
@@ 5 Discussion and Conclusion
Besides the benefits shown in experiments, our method possesses additional advantages worth discussing.
Cost-effective.
Our method is highly cost-effective. In our experiments, we needed to annotate an average of 3,000 texts, each approximately 60 words in length for each experiment. Using the standard practice of cross-validation, hiring two workers on Amazon Mechanical Turk (MTurk) would result in a conservative cost estimate of \$$\$350$ per worker, plus a \$$\$100$ MTurk fee, totaling \$$\$850$. In contrast, our approach incurs a total cost of approximately \$$\$183$, which includes \$$\$2$ for the embedding model, \$$\$92$ for the weak models, and \$$\$79$ for the advanced model. This means our method costs at most $21.5\%$ of the total price of hiring human workers, not to mention the additional computational cost of training.
Extensible.
Our method is extensible, allowing researchers to adapt or replace the algorithms within our framework to fit specific research needs. For instance, while we used UMAP, K-means, and MapReduce in our experiments for exemplar selection and prompt generation, these algorithms are not fixed and can be easily substituted with alternatives. This flexibility enables the method to be tailored to different types of data and research objectives, making it a versatile tool for a broad range of political science applications. Additionally, our approach is designed to integrate the latest LLM advancements seamlessly, ensuring that it remains at the forefront of technological progress without requiring significant adjustments or incurring additional costs. For example, if OpenAI releases GPT-5 in the future, our framework supports users to add this new model to their research just by modifying the configure file. This adaptability future-proofs our method, allowing it to consistently deliver cutting-edge performance as LLM technology evolves.
Limitation.
Our method has several limitations. First, it is currently limited to categorical classification and is not suitable for tasks that require measuring intensity, degrees, or rankings. For example, our approach lacks the capability to accurately assess political ideology on a traditional 5-point scale, such as distinguishing between strong Democrats, moderate Democrats, Independents, Republicans, and strong Republicans. In future research, we plan to extend our method to handle pairwise comparisons, enabling us to better capture the intensity of party identification and other continuous variables.
Second, our method is currently limited to text data and cannot be applied to more complex data types such as images, videos, and audio, all of which are increasingly gaining attention in political science (e.g., Torres_Cantú_2022; Girbau_Kobayashi_Renoust_Matsui_Satoh_2024; Torres_2024). Expanding our method to handle these diverse forms of data is a crucial direction for future research, allowing us to better address the growing variety of data sources in the field.

### v2

@@ 4.1 Sentiment Analysis and Stance Detection toward Brett Kavanaugh’s SCOTUS Nomination Using Twitter Data
Figure 5: Measuring Sentiment toward Kavanaugh: Comparison of F1 Scores across Different Methods
Note: The gray bars represent the classification results using LLMs directly, without any enhancement from our method. The green and orange bars indicate the results after applying our three-stage method, demonstrating the effectiveness of each stage in improving classification performance.
Following sentiment classification, we extended our method to stance detection. Bestvater and Monroe’s study evaluated multiple text classifiers for stance detection and concluded that models or dictionaries initially designed for sentiment analysis often perform poorly on stance tasks, suggesting a need for models specifically trained for stance recognition. These findings are summarized in Table D0. Contrary to Bestvater and Monroe’s recommendation that a newly trained model is essential for accurate stance classification, our method demonstrates that reliable classification can be achieved by simply changing and optimizing prompts.
@@
As in the previous analysis, we began with the unoptimized heuristic prompt shown in Figure 6. Using this unoptimized prompt, both Mistral-medium-latest and GPT-3.5-turbo achieved scores below 60%, highlighting their difficulty in distinguishing between sentiment and stance. For example, both models incorrectly labeled the following tweet as “oppose”:
@@
”RT @atensnut Democrats can’t just ’move on’ and jump on the bandwagon of sketchy allegations against Kavanaugh without accepting the egregiousness of turning their backs on the victims of Bill Clinton.”
@@
The LLMs mistakenly categorized this as “oppose” because the tweet carries a tone of negativity and distrust, implicitly criticizing the Democratic Party for supporting the allegations against Kavanaugh while highlighting perceived hypocrisy. This example illustrates how LLMs, when prompted with a simple heuristic, can conflate sentiment with stance—misinterpreting negative sentiment toward a third party as opposition to the main subject. It underscores the importance of using structured, task-specific prompts that are designed to capture the nuanced distinctions required for accurate stance detection.
We then provided LLMs with enhanced prompts produced by our automatic prompt optimization method (Figure 6), offering clear and structured guidance that significantly improved their classification accuracy. Notably, the enhanced prompt goes beyond simple factual classification; for instance, it correctly identifies tweets as “approve” when they present evidence undermining Kavanaugh’s accusers. It also effectively captures and summarizes emotional cues that may imply support for Kavanaugh—such as mocking his opponents or expressing frustration at those obstructing his confirmation—nuances that are often misclassified as opposition when using heuristic-based prompting. These cases highlight how negative tone or sentiment alone can misleadingly suggest stance, underscoring the importance of prompts that distinguish between emotional expression and political position.
@@
This refined prompt ensures both factual and emotional content are considered, leading to more accurate and nuanced classifications. We further validated the enhanced prompt against the definitions established in Figure 3 and found strong alignment, reinforcing the conceptual distinction between sentiment and stance. The enhanced prompt yielded substantial performance improvements: Mistral-medium and GPT-3.5 achieved F1 scores of 91.69% and 92.59%, respectively—representing a ~36% increase over the initial zero-shot performance using a simple heuristic prompt, as shown in Figure 7. In the four-shot [^11] Detailed few-shot prompts are provided in Appendix D2. prompting scenario, the enhanced prompt enabled these models to achieve over a 20% improvement, further demonstrating its value in producing accurate classifications. [^12] The detailed resulting performance metrics for four-shot prompting are presented in Appendix D2. Notably, by incorporating chain-of-thought reasoning and a final validation step using the Mistral-large-latest model as a judge, our approach surpassed traditional supervised classifiers, with F1 scores exceeding 95%.
@@
Figure 6: Zero-Shot Prompts for Analyzing Tweets about Brett Kavanaugh
@@
Note: Italicized sections reflect additions made after applying automatic prompt optimization. The optimized prompt succinctly outlines conditions under which text should be classified as “oppose” or “support.”
Our method demonstrates superior performance compared to unoptimized LLM predictions, largely due to the optimized prompt generated in the second stage of our approach. This shows an important consideration for political scientists: when using LLMs for text classification, relying on unoptimized labeling alone can lead to significant misclassifications, as LLMs may struggle with complex concepts like sentiment and stance. However, by providing a few exemplars and instructing the LLM to generate a more refined prompt with simple, targeted guidelines, we achieve outstanding results. Furthermore, our method outperforms both dictionary-based and traditional supervised learning approaches, all without the need to train a new model—contrary to the recommendation by Bestvater and Monroe (2023) for improved classification. We successfully minimized the cost associated with model training while maximizing classification accuracy.
Figure 7: Stance Detection toward Kavanaugh: Comparison of F1 Scores across Different Methods
We further assessed the relationship between sentiment and stance in our results by coding sentiment (positive as 1, negative as 0) and stance (support as 1, oppose as 0), using both Pearson correlation and the Jaccard Index to capture alignment between these two dimensions. [^13] The Jaccard Index, which measures similarity by dividing the size of the intersection by the size of the union between sets, indicates greater overlap when higher, suggesting potential confusion in classifications between sentiment and stance (levandowsky1971distance). Detailed formulas and interpretation for both metrics are provided in  D. For zero-shot results, the unoptimized application of LLMs showed a Pearson correlation of 0.38, whereas the optimized model returned a considerably lower correlation of 0.14, indicating weaker association between sentiment and stance. Similarly, the Jaccard Index for unoptimized zero-shot was 0.349, while optimization reduced it to 0.28, reinforcing that our method more effectively distinguishes stance from sentiment. A similar pattern emerged in the four-shot results: the unoptimized correlation was 0.31 compared to 0.07 in the optimized version, and the Jaccard Index dropped from 0.374 to 0.20. A statistical test confirmed that the difference in correlation coefficients was significant ($p<0.05$), demonstrating that the optimized approach more effectively separates the two constructs.
@@ 4.2 Classifying Campaign Ads Tones in the 2018 Election
The medium through which political communication is delivered plays a critical role in shaping the message’s tone and its audience reach. In a recent study, fowler2021political examined the impact of Facebook as a medium on the tone of political advertisements. They proposed that ads on Facebook are more likely to adopt a negative tone compared to other platforms [^14] For a broader discussion on the prevalence and dynamics of negative campaign tone across various media, see Lau and Rovner’s review article ”Negative Campaigning” in the Annual Review of Political Science (lau2009negative)..
@@
We applied the same three-stage approach as in sentiment analysis and stance detection, this time focusing solely on reporting consensus mechanism results of the Chain-of-Thought (CoT) and Judge prompting methods. We evaluated their performance under both zero-shot and few-shot prompting conditions [^16] Detailed unoptimized and optimized results are available in the Appendix.. The final F1 scores, as shown in Table 1, indicate notably low performance in the ”attack” and ”contrast” categories. Specifically, the average F1 score for ”attack” is approximately 55%, while for ”contrast,” it is even lower, averaging around 50%. It is important to note that these F1 scores were calculated against gold-standard human labels, which are particularly noisy in this dataset, potentially contributing to the reduced accuracy.
@@
Class 0-Shot CoT 0-Shot Judge 6-Shot CoT 6-Shot Judge
promote 92.85% 93.23% 92.60% 93.14%
contrast 53.85% 45.67% 51.80% 49.78%
attack 57.19% 53.68% 56.48% 58.56%
@@
These findings facilitated a re-examination of the data, particularly in cases where labels generated by multiple LLMs differ, mimicking a scenario with multiple human coders who occasionally disagree. Such discrepancies can highlight conceptual confusion or subtle ambiguities in the text, which may lead to controversial or inconsistent labeling decisions. Examples of such label inconsistencies, as shown in Table 2, highlighting the need to question the consistency and reliability of the so-called gold standard human labels. In the first text, which was labeled as ”contrast” by human coders, the primary focus is on criticizing career politicians for failing to prevent the destruction of industries, without directly contrasting specific policies with those proposed by others. On the contrary, the second piece, labeled as ”promote” by human coders, explicitly criticizes the Republican Party in the state legislature while calling for support for other candidates. This message could reasonably be interpreted as both a promotion of alternatives and a contrast with the criticized party. Despite the similarities between these two ads—both of which criticize opponents and call for the support of others—the human labels differed. On the other hand, GPT and the final chain-of-thought method remained consistent in their classifications, demonstrating a more stable and coherent approach to labeling.
@@
Notes: The left panel of this figure shows the effect of medium on campaign ad tones across different labeling approaches, including initial human-labeled data before removing controversial labels
Given the noise within the human-labeled data, we proceeded to remove all controversial observations—those instances where two LLMs produced conflicting labels. Table 3 highlights the substantial improvements in F1 scores across all categories following this adjustment. Notably, the contrast and attack categories saw over a 30% increase in F1 scores, demonstrating how filtering ambiguous cases can enhance label quality and model evaluation.
After this refinement, we reran the downstream fixed-effects model using outcomes exclusively from the optimized LLM classification on both matched data. The yellow line in the right panel of Figure 8 illustrates this estimator across each category. Additionally, we applied the fixed-effects model to this matched sample using human-labeled outcomes, shown by the blue lines in the same panel. The black line continues to represent estimates derived from human labels across the full sample. Notably, we observed that the estimates produced by LLM labeling on the matched sample closely align with those from human labeling on the matched sample. This similarity between LLM and human labeling on non-controversial data indicates the reliability of optimized LLM classifications in producing consistent results. Moreover, when comparing estimates derived from human labeling on the full sample of only 3,000 campaign ads to those based only on the matched sample, we observed notable differences, particularly for the Contrast category. This finding indicates that even within a full sample of 3,000 human-labeled ads, labeling inconsistencies can introduce noise that affects downstream analyses. Consequently, ensuring consistency in human labeling is crucial for achieving reliable estimates and minimizing potential biases in subsequent analysis.
@@
Additionally, Table 1 revealed that zero-shot metrics surpassed few-shot performance. This anomaly can be attributed to the sampling of 80 examples for few-shot prompting during the dynamic exemplar selection stage. Given the noise within the human labels, these samples might have included incorrectly labeled examples, impacting the accuracy of few-shot promptings. This reliance on potentially flawed data could degrade the performance of few-shot prompts to levels below that of zero-shot prompts, which remain uninfluenced by human input.
This experiment reveals a significant advantage of our method: its capacity to identify and flag inconsistencies in human labeling. Although human-coded data is often considered the “gold standard,” our findings reveal that this assumption can be problematic, particularly when dealing with contentious classifications. Human-coded classifications of ad tone are especially prone to noise due to the subjective nature of tone interpretation and the nuanced distinctions between categories like ”attack” and ”contrast.” These categories often involve overlapping criteria, with small variations in content or emphasis leading to disagreements among coders. Additionally, individual biases, contextual interpretations, and a lack of universally agreed-upon definitions exacerbate inconsistencies in human annotations. [^18] Inconsistencies among human coders are well-documented in the literature; see, for example, lombard2002content, neuendorf2002content, and artstein2008inter. For ad tone classifications, the overlap between categories such as ”attack” and ”contrast” can further complicate consistent labeling, as highlighted in studies of political communication.This analysis demonstrates that absolute reliance on human coding can be risky, especially when only a single coder is used. By detecting and flagging these discrepancies, our method enhances the robustness of classification results and allows researchers to address biases inherent in human-coded data before they affect downstream interpretations.
@@
Table 3: Analyzing Ads Tones: F1 Scores after Addressing Discrepancies: CoT Prompting vs. Judge Prompting (Zero-Shot)
@@ 5 Discussion and Conclusion
In this paper, we introduce a novel three-stage LLM-based framework for text classification specifically tailored to political science research. Our approach addresses the challenges of traditional supervised learning and language models, which often require labor-intensive and computationally expensive feature engineering and fine-tuning. By leveraging inherent capabilities of in-context learning provided by LLMs, our method eliminates the need for such manual interventions.
Crucially, our framework offers an innovative solution to one of the biggest challenges in using LLMs for text classification: the heavy reliance on prompt quality to achieve accurate outputs. Through automatic prompt optimization, our method allows LLMs to generate task-specific prompts that are precisely tailored to the task at hand, ensuring that models accurately interpret and classify the text. This optimization greatly improves the accuracy and reliability of classification results, minimizing the need for extensive human intervention in prompt design.
Additionally, our approach incorporates dynamic exemplar selection, which resolves the issue of irrelevant exemplar selection in few-shot learning. By dynamically choosing exemplars most relevant to the query text, our method prevents suboptimal classification results, further enhancing the overall performance and adaptability of the LLMs in diverse text classification tasks.
Finally, to ensure robustness and interpretability, our framework integrates a consensus mechanism in which multiple LLMs are used to cross-validate predictions. When discrepancies arise, a judging LLM evaluates the reasoning behind conflicting outputs to determine the most accurate classification. This structured adjudication process mirrors human multi-coder workflows and significantly reduces noise and uncertainty in model predictions, providing greater confidence in the final results.
Our method offers significant extensibility, making it a highly adaptable tool for political science research. One key advantage is its modular design, allowing researchers to easily customize or replace the algorithms within our framework to suit specific research needs. For example, although we utilized UMAP, k-means, and MapReduce for exemplar selection and prompt generation in our experiments, these components are not fixed; they can be replaced with alternative algorithms tailored to the type of data or research objectives at hand. This flexibility makes the framework applicable to a wide variety of political science tasks and data types.
Moreover, our approach is designed to seamlessly integrate with the latest advancements in LLM technology, ensuring it remains cutting-edge without the need for significant modifications or added costs. For instance, if a newer model like GPT-x were to be released, researchers could easily incorporate it into our framework without any coding adjustments. This adaptability effectively future-proofs the method, enabling it to continually leverage state-of-the-art LLM capabilities. As long as LLMs are used, our method will remain applicable and relevant, providing a sustainable and up-to-date tool for political science research.
Last but not least, our method offers a cost-effective framework that achieves high accuracy through a consensus mechanism, making it a valuable tool for researchers seeking reliable yet accessible classification solutions. [^19] Table A4 presents a cost and performance comparison between GPT-4-turbo and GPT-3.5-turbo. Our method, which relies on a consensus mechanism using GPT-3.5-turbo calls, achieves similar accuracy to GPT-4-turbo while reducing costs by more than a factor of 5 with three stages combined.
Future Directions.
While our method demonstrates strong performance in categorical text classification, it also opens up several avenues for future research. One promising direction involves extending the framework to tasks that require measuring intensity, degrees, or rankings—such as estimating political ideology on a continuous scale. For example, distinguishing between strong Democrats, moderate Democrats, Independents, Republicans, and strong Republicans on a 5-point Likert scale remains a challenge. This is due in part to the fact that large language models (LLMs) do not natively encode ordinal relationships during pre-training. Although they are highly effective at generating and classifying categorical language patterns, they lack a built-in understanding of ordinal or continuous gradations—a limitation that has long been recognized in the field of text scaling (see grimmer2013text; lowe2013validating). Developing methods that integrate ordinal awareness or continuous scaling into prompt-based classification could be especially impactful for political science applications.
Another important future direction is the extension of this framework beyond text to support multimodal data types such as images, videos, and audio, which are increasingly central to political communication and behavior research (e.g., Torres_Cantú_2022; Girbau_Kobayashi_Renoust_Matsui_Satoh_2024; Torres_2024). Adapting the principles of automatic prompt optimization, dynamic exemplar selection, and LLM-based consensus to handle these diverse data forms would expand the utility of our approach and enable researchers to analyze a broader range of political content in real-world contexts.